vale1213:

Given: If CR = 4, RA = 6, and RS = 10, find AB. 6 2/3 15 16 2/3 25

darkknight:

Is there a drawing or anything of the sort to help us with this problem?

darkknight:

You can click the attach file button down below.

vale1213:

no there isn't because it wont let me bring the drawing into the question.

vale1213:

and ok

darkknight:

If that doesn't work, there is also a draw feature where you can manually draw the picture.

darkknight:

Yes, you have to click attach file from here.

vale1213:

alright

vale1213:



vale1213:

the black triangle was supposed to be an A

darkknight:

Okay, np.

darkknight:

So 

darkknight:

now we have a ratio that we can set up. So CR/RS = CA/AB. CR is 4, and RS is 10. CA is 6+4 or 10, and we don't know AB.

darkknight:

So that means that 4/10 = 10/AB, can you solve for AB from here?

vale1213:

yea i think i can figure it out thank you

darkknight:

np, have good day, also be sure to click best response (up to you though). :)

vale1213:

i really appericaite it and sure thing :)

YNWMelly:

:)